Find the p-value for the hypothesis test. A random sample of size 51 is taken. The sample has a mean of 423 and a standard deviation of 84.

H0: µ = 400

Ha: µ > 400

The p-value for the hypothesis test is .

Your answer should be rounded to 4 decimal places.
Answer #1

This is the right tailed test .

The null and alternative hypothesis is ,

H0 :  \mu =400

Ha : \mu >400

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (423 -400) / 84 / \sqrt 51

= 2.04

P-value = 0.0207